# Context Matters: Pushing the Boundaries of Open-Ended Answer Generation with Graph-Structured Knowledge Context

### v2

@@ Abstract
In the continuously advancing AI landscape, crafting context-rich and meaningful responses via Large Language Models (LLMs) is essential. Researchers are becoming more aware of the challenges that LLMs with fewer parameters encounter when trying to provide suitable answers to open-ended questions. To address these hurdles, the integration of cutting-edge strategies, augmentation of rich external domain knowledge to LLMs, offers significant improvements. This paper introduces a novel framework that combines graph-driven context retrieval in conjunction to knowledge graphs based enhancement, honing the proficiency of LLMs, especially in domain specific community question answering platforms like AskUbuntu, Unix, and ServerFault. We conduct experiments on various LLMs with different parameter sizes to evaluate their ability to ground knowledge and determine factual accuracy in answers to open-ended questions. Our methodology GraphContextGen consistently outperforms dominant text-based retrieval systems, demonstrating its robustness and adaptability to a larger number of use cases. This advancement highlights the importance of pairing context rich data retrieval with LLMs, offering a renewed approach to knowledge sourcing and generation in AI systems. We also show that, due to rich contextual data retrieval, the crucial entities, along with the generated answer, remain factually coherent with the gold answer.
Figure 1: Schematic of the GraphContextGen framework.
@@ 1 Introduction
In the field of artificial intelligence, LLMs Roberts et al. (2020); Kaplan et al. (2020) stand out as significant breakthroughs. They highlight the advanced capabilities of machines in both text understanding Lian et al. (2023) and generating Wei et al. (2023). While LLMs usually produce high-quality outputs they often struggle in areas with limited resources Chen et al. (2023); Guu et al. (2020b). They also have limitations due to knowledge cutoff dates. In many cases, they hallucinate McKenna et al. (2023) when faced with very specific, low-resource or domain specific topics. There is also a trade-off between quality and the billions of parameters Xu et al. (2023), which makes the LLM challenging to explore and analyze for researchers with limited resource capacity. Recently, to address these challenges, grounding of LLMs  [^1] https://techcommunity.microsoft.com/t5/fasttrack-for-azure/grounding-llms/ba-p/3843857 and the Retrieval-Augmented Generation (RAG) method Yu et al. (2023) have been proposed. This approach allows large models to actively utilize the information from external databases as they operate. This feature is particularly useful as it means the models can provide answers that are not just based on their original training data, but are also detailed, current, and well-informed by accessing a wider range of information sources. This opens up the potential for even more factually accurate and contextually relevant responses in real-time.
A pivotal aspect shaping the effectiveness of knowledge grounding is its choice of retrieval Lewis et al. (2021) mechanism. Even though traditional text-based retrieval methods were favored for pulling content from diverse sources, the increasing complexity of questions and the demand for more than simple keyword matching is prompting a change in the research landscape. Moreover, prevailing state-of-the-art (SOTA) methodologies in text-based retrieval grapple with determining the optimal chunk size  [^2] https://www.pinecone.io/learn/chunking-strategies/, particularly when indexing and querying. This inconsistency in chunk delineation can introduce variability in results, especially when interfacing with diverse datasets that exhibit heterogeneous characteristics in terms of content density, topic distribution, and structural composition. Graph-based retrieval systems are gaining traction in this regard. Instead of just casting a wide net, they delve into structured data where interconnected nodes and edges encapsulate intricate relationships, ensuring a deeper semantic understanding. This structured approach not only brings forth more contextually pertinent results by cutting down the ambiguity often seen in keyword-centric searches Zhang et al. (2021), but it is also adept at swiftly navigating complex queries by traversing interconnected knowledge. The dynamic nature of these systems further allows them to adapt to ever-evolving data, paving the way for unearthing unexpected insights and forming indirect, transitive connections among diverse entities.
@@
This grounding technique is vital in various text generation sectors, including dialogue systems Li et al. (2022), open-domain question answering Lu et al. (2023), and novelty-controlled paraphrasing Xie et al. (2023). StackOverflow’s recent initiative, OverflowAI  [^3] https://stackoverflow.blog/2023/07/27/announcing-overflowai/, focuses on refining search, adding new features, enhancing code, and fostering discussion platforms. Simultaneously, the trend in automated answer generation on community question answering platforms promises timely, accurate information delivery, reducing errors, and ensuring immediate knowledge access. While past research has centered on selecting answers from prior questions or summarizing them, our study uniquely harnesses LLMs to craft answers specifically for these platforms.
@@
Therefore, we introduce the  GraphContextGen framework that incorporates graph-based retrieval and offers enhanced context to LLMs. Our approach centers on ensuring the factual accuracy of the generated answers through knowledge graph-based context enhancement. In addition, we carry out extensive experiments with a wide range of LLMs in terms of different parameters for answer generation. We primarily target low-resource domains like AskUbuntu  [^4] https://askubuntu.com/, Unix  [^5] https://community.unix.com/, and ServerFault  [^6] https://serverfault.com/, aiming to gauge the effectiveness of LLM-generated answers in these specialized areas and evaluate their resilience against catastrophic forgetting.
Contribution: The key contribution of this paper is as follows.
@@
  • We introduce a framework named GraphContextGen which integrates a graph-based retrieval mechanism in conjunction with enhancement of context using knowledge graph for CQA answer generation. This combined approach consistently outperforms previous state-of-the-art (SOTA) methods. Moreover, employing the instruction tuning strategy with LLMs further enhances performance (see Section 4).
@@
  • We perform a detailed retrospective analysis comparing actual answers with those generated by our framework. This analysis aims to understand the factual alignment between generated and actual answers, noting that generated answers typically align factually with the actual ones (see Figure 3).
@@ 2 Related Work
Large language models: Recent progress in LLMs Naveed et al. (2023); Douglas (2023) has positioned them as dominant successors to earlier transformer-based models for tackling generation tasks in natural language. Over the past several years, there has been a surge in introducing LLMs  [^7] https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard like GPT3 Brown et al. (2020), Llama2 Touvron et al. (2023a) etc. for this purpose. The newly introduced LLMs have demonstrated the ability to grasp and interpret context, subsequently producing human like responses. These models generate fluent and high quality answers very well for general questions Yang et al. (2023c). So, these models are often used as the base architecture for various generation based tasks such as long Q&A Saad-Falcon et al. (2023); Kamalloo et al. (2023), multiple choice Q&A Robinson et al. (2023), dialogue systems Deng et al. (2023); Hudeček and Dušek (2023); Valvoda et al. (2022); Snell et al. (2022), chatbot systems Lee et al. (2023a); Wei et al. (2023). However, despite the high capability of these methods, they generate factually incorrect Shuster et al. (2021) responses and are oftentimes not aligned with the actual question for domain specific knowledge. Generally, this issue of LLMs is called hallucination Dhuliawala et al. (2023); Mündler et al. (2023). Due to the cutoff knowledge, models sometime fail to provide faithful answer for a general question. To tackle this problem, recent studies Shuster et al. (2021) proposed a method of augmenting retrieved knowledge from diverse sources as a context to LLMs.
Retrieval augmented generation: Retrieval-augmented generation (RAG) Manathunga and Illangasekara (2023); Ram et al. (2023) is a framework that merges the capabilities of large pretrained language models with the advantages of external retrieval or search mechanisms. Earlier, augmenting language models with retrieved knowledge had shown effective performance for knowledge-intensive tasks Guu et al. (2020a). It is particularly useful in scenarios where a language model needs to pull in specific factual information from an extensive corpus to generate relevant and accurate outputs. In recent studies, the researchers attempt to augment external knowledge through knowledge graph prompting Wang et al. (2023), few shot domain adaptation Krishna (2023), retrieval-generation synergy Shao et al. (2023), knowledge graph-based subgraph retrieval augmented generation Kang et al. (2023), augmented adapter retriever (AAR) Yu et al. (2023).
Instruction tuning: Although LLMs have already shown impressive generation capabilities, oftentimes they cannot align the user’s objective with its training objective Zhang et al. (2023). This is because while in general LLMs are trained to minimize the contextual word prediction error, users want to receive responses given some instructions of their choice. To mitigate that, various methods such as supervised fine tuning (instruction tuning) von Werra et al. (2020), RLHF Ziegler et al. (2019); Stiennon et al. (2020); Ouyang et al. (2022); Gao et al. (2022), RLAIF Lee et al. (2023b) have been used to enhance the capabilities and controllability of LLMs. Also, parameter efficient techniques Mangrulkar et al. (2022) have been proposed to fine tune less number of model parameters and more adapter’s parameters.
@@
Answer generation: Over the years several approaches such as feature based methods Wang et al. (2009); Wang and Manning (2010), convolutional neural networks (CNNn) Severyn and Moschitti (2015); Rao et al. (2017), recurrent neural neteworks (RNNn) Wang and Nyberg (2015), attention mechanism Tan et al. (2016); dos Santos et al. (2016) have been proposed for answer selection and summarization. Some studies have leveraged additional information to balance the information between questions and answers. For example, user models Wen et al. (2018); Li et al. (2017), latent topics Yoon et al. (2018), external knowledge Shen et al. (2018), or question subjects Wu et al. (2018) have been utilized to compensate for this information imbalance. Several answer summarization techniques are proposed for low-resource domains, such as condensing technical answers into concise summaries. Some recent research focuses on summarizing diverse content on StackOverflow Chengran et al. (2022), using AnswerBot–an answer summary generator Xu et al. (2017), Opiner–which summarizes API reviews Uddin and Khomh (2017), extracting key sentences to guide developers on StackOverflow Nadi and Treude (2019), and multi-document summarization Xu and Lapata (2020). In the realm of answer summarization, numerous studies Ganesan et al. (2010); Naghshzan et al. (2021) harness graphical structures, leverage existing graph-based summarizers Mihalcea and Tarau (2004); Erkan and Radev (2004); Kazemi et al. (2020), and employ graph-centric measures. In the age of LLMs, some research has centered around controlled summary generation via effective keyword-based prompting He et al. (2022).
@@ 3 Dataset
In this experiment, we select three domain-specific datasets from open-source CQA platforms: AskUbuntu, ServerFault, and Unix, all of which originate from a low-resource domain with minimal properly annotated data available on these topics. These datasets, considered from June 2023, includes questions (title and body), a list of answers, an accepted answer flag, tags for the questions, and the posting dates and times for both questions and answers. For each question, the accepted answer serves as the ground truth. Due to the limited resources of the datasets and the high expenses associated with human involvement, we opt not to use human annotations. Further we apply several filtering procedures, such as duplicate question removal, non-specific answer removal, and length constraints (token limit in LLM) resulting in our dataset. Adopting the temporal splitting approach inspired by Hazra et al. (2021, 2023), we consider training set from 2019-2020 and test dataset from 2021-2023. Due to resource limitations, we randomly sample test dataset, the details of which are provided in Table 1. From the training set of each dataset, we construct an instruction-tuning dataset by pairing questions and answers in the format `$[INST]Question[\textbackslash INST]Answer:actual\_answer^{\prime}$. We prepare this instruction data for each dataset.
@@ 5 Experimental setup
  1. (i)
    Pre-LLM era baselines: We compare our approach with SOTA answer generation/summarization works such as AnswerBot Xu et al. (2017); Cai et al. (2019), GenQA Hsu et al. (2021) and TechSumBot Yang et al. (2023a) (see Appendix A.2). Due to unavailability of the codebase and unclear implementation details, we could not compare this paper Deng et al. (2019) with our method.
  2. (ii)
    Zeroshot LLMs: In this setting, we use various competitive LLMs to generate the answer of the given question. The LLMs are of different parameter sizes (7B to 40B). Such a choice enables us to understand how well models with diverse parameter sizes perform in zero shot setting.
  3. (iii)
    [w/o INST] TextGen: In this setup, we use a vector database (chromaDB [^10] https://docs.trychroma.com/getting-started and FAISS [^11] https://python.langchain.com/docs/integrations/vectorstores/faiss) containing all the training set questions. We compute contextual similarity between query $q$ and all the questions in database. We rank the questions in database based on the cosine similarity scores (higher scores get top ranks) and retrieve top $k$ questions. Further we use the top $k$ questions and their actual answers as few shot examples to the pretrained LLM for generating the answer.
  4. (iv)
    [w/o INST] TextContextGen: In this setup, we retrieve the top $k$ questions and their answers using the same method as  [w/o INST] TextGen. Subsequently, we use context enhancement component of our approach to enhance the context. Further we provide the enhanced context and the query as input to the pretrained LLM and obtain the generated answer.
  5. (v)
    [w/o INST] GraphGen: In this setup, we use the Retriever module of our algorithm to retrieve the top $k$ questions. We use $k$ questions and their answers as few shot examples to the pretrained LLM for generating the answer.
  6. (vi)
    [w/o INST] GraphContextGen: We follow the retrieval step from  [w/o INST] GraphGen. Further we use our ContextEnhancer module to enhance the context.
  7. (vii)
    FineTuned Gen Zero-Shot: In this setting, we use instruction fine tuned LLM in zero shot settings. Here, we pass the questions as input and the fine tuned LLM generates the answer.
  8. (viii)
    TextGen: This setup is same as [w/o INST] TextGen. However, we use our instruction fine tuned LLM for generation.
  9. (ix)
    GraphGen: This setup is same as [w/o INST] GraphGen. However, we use our instruction fine tuned LLM for generation.
@@ 6 Results
Zero-shot answer generation by different LLMs: For the AskUbuntu dataset, the Llama2 (13B) model consistently achieves the top BERTScore (0.810) compared to the other existing models. For ROUGE 1, ROUGE L metrics also, this model leads with scores of 0.227 and 0.211, respectively. For the Unix dataset, the Llama2 (13B) model remains dominant, particularly in BERTScore (0.806). Vicuna (13B) is best in terms of ROUGE 1 (0.181), while Llama2 (13B) is best in terms of ROUGE L (0.175). For the ServerFault dataset, the trend remains similar: Llama2 (13B) dominates with a BERTScore of 0.819; Vicuna (13B) leads in terms of ROUGE 1 (0.185) and Llama2 (13B) in terms of ROUGE L (0.176).
We observe that the performances are not always directly proportional to the model size. Llama2 (13B) most often outperforms bigger models like Gpt-neox (20B) and Falcon (40B). Similarly, models of comparable sizes, such as Falcon (7B), MPT (7B), and StackLlama (7B), display varied performances, emphasizing that model architecture and training methods are equally crucial as is the size.
Main results: Table 3 shows the baseline results and compares them with our proposed method. In the Pre-LLM era, AnswerBot showcases competitive results with BERTScore of 0.803, 0.791, and 0.802 for AskUbuntu, Unix, and ServerFault, respectively. GenQA tends to underperform with the lowest BERTScore of 0.55 for the Unix platform. Among the methods from this era, AnswerBot also achieves the highest FactSumm score across all platforms.
In the results pertaining to the LLM era, the reference model for generating answers is always set to Llama2 (13B) which performed the best in the zero-shot setting [^16] Interestingly, the results remain unchanged even if we use Llama2 (70B).. For all the three datasets, our models GraphContextGen outperforms all other baselines in terms of BERTScore. In terms of FactSumm, GraphContextGen is best for two out the three datasets (AskUbuntu and Unix). This indicates that our model produces factually more correct answers compared to the different baselines. For the ROUGE 1 and ROUGE L once again GraphContextGen shows competitive performance in two out of the three datasets (Unix and ServerFault). In general, the models of the LLM era are always better in terms of these two metrics compared to the models of the pre-LLM era.
Grounding of the generated answers: We use the UniversalNER Zhou et al. (2023) model to identify the entities present in the ground truth answer and the generated answer. The Jaccard similarity between the entity sets in the ground truth and the generated answers for our model are 0.85, 0.75 and 0.79 for the Askubuntu, Unix, and ServerFault datasets respectively (see Figure 3 (A)). The overlap in the number of triplets for the three datasets is shown in Figure 3 (B). The numeric values in both the figures point that the answers generated by our model are rich in entity and relationships that are actually present in the ground truth answer.

### v3

@@ Abstract
This paper introduces a novel framework that combines graph-driven context retrieval in conjunction to knowledge graphs based enhancement, honing the proficiency of LLMs, especially in domain specific community question answering platforms like AskUbuntu, Unix, and ServerFault. We conduct experiments on various LLMs with different parameter sizes to evaluate their ability to ground knowledge and determine factual accuracy in answers to open-ended questions. Our methodology GraphContextGen consistently outperforms dominant text-based retrieval systems, demonstrating its robustness and adaptability to a larger number of use cases. This advancement highlights the importance of pairing context rich data retrieval with LLMs, offering a renewed approach to knowledge sourcing and generation in AI systems. We also show that, due to rich contextual data retrieval, the crucial entities, along with the generated answer, remain factually coherent with the gold answer.
@@ 1 Introduction
In artificial intelligence, Large Language Models (LLMs)Roberts et al. (2020); Kaplan et al. (2020) have revolutionized text understandingLian et al. (2023) and generation Wei et al. (2023). Despite their impressive capabilities, LLMs struggle in low-resource settings Chen et al. (2023); Guu et al. (2020), are constrained by knowledge cutoffs, and often produce hallucinations McKenna et al. (2023). Additionally, managing the trade-off between quality and the vast number of parameters Xu et al. (2023) presents challenges, particularly for researchers with limited resources.
Figure 1: GraphContextGen framework.
To overcome these limitations, new methods such as grounding LLMs  [^1] https://techcommunity.microsoft.com/t5/fasttrack-for-azure/grounding-llms/ba-p/3843857 and Retrieval-Augmented Generation (RAG) Yu et al. (2023) have been proposed. These approaches enable models to access external databases, enhancing their responses with current, detailed, and accurate information.
A critical aspect of effective knowledge grounding is the retrieval mechanism Lewis et al. (2021). Traditional text-based retrieval methods are evolving to handle more complex questions, moving beyond simple keyword matching. Current techniques often struggle with determining optimal chunk sizes  [^2] https://www.pinecone.io/learn/chunking-strategies/ for indexing and querying, leading to inconsistent results. Graph-based retrieval systems offer a solution by capturing intricate relationships through structured data, providing deeper semantic understanding and more contextually relevant results Zhang et al. (2021). These systems adapt to evolving data, uncovering insights and forming connections among diverse entities.
@@
This technique is vital in various applications, including dialogue systems Li et al. (2022), open-domain question answering Lu et al. (2023), and novelty-controlled paraphrasing Xie et al. (2023). For example, StackOverflow’s OverflowAI  [^3] https://stackoverflow.blog/2023/07/27/announcing-overflowai/ aims to refine search and enhance code and discussion platforms. Automated answer generation on community Q&A platforms promises timely and accurate information, reducing errors and providing immediate knowledge access. Unlike past research, our study uniquely employs LLMs to generate tailored answers for these platforms.
@@
We introduce the GraphContextGen framework, which combines graph-based retrieval with LLMs to enhance context and ensure factual accuracy. Our extensive experiments in low-resource domains like AskUbuntu [^4] https://askubuntu.com/, Unix  [^5] https://community.unix.com/, and ServerFault  [^6] https://serverfault.com/ demonstrate the effectiveness and resilience of LLM-generated answers, even in specialized areas.
Contribution: The key contribution of this paper is as follows. {stylishframe}
@@
  • We introduce GraphContextGen, a framework that integrates graph-based retrieval with context enhancement using a knowledge graph for CQA answer generation. This approach consistently outperforms previous SOTA methods. Additionally, instruction tuning with LLMs further improves performance (see Section 4).
@@
  • We conduct a detailed retrospective analysis to compare actual answers with those generated by our framework, focusing on their factual alignment. The generated answers typically align well with the actual ones (see Figure 2).
@@ 2 Related Work
Over the years several approaches such as feature based methods Wang et al. (2009); Wang and Manning (2010), CNN Severyn and Moschitti (2015); Rao et al. (2017), RNN Wang and Nyberg (2015), attention mechanism Tan et al. (2016); dos Santos et al. (2016) have been proposed for answer selection and summarization. Some recent research focuses on summarizing diverse content on StackOverflow Chengran et al. (2022), using AnswerBot–an answer summary generator Xu et al. (2017), Opiner–which summarizes API reviews Uddin and Khomh (2017), extracting key sentences to guide developers on StackOverflow Nadi and Treude (2019), and multi-document summarization Xu and Lapata (2020). In the realm of answer summarization, numerous studies Ganesan et al. (2010); Naghshzan et al. (2021) harness graphical structures, leverage existing graph-based summarizers Mihalcea and Tarau (2004); Erkan and Radev (2004); Kazemi et al. (2020), and employ graph-centric measures. In the age of LLMs, some research has centered around controlled summary generation via effective keyword-based prompting He et al. (2022).
@@ 3 Dataset
In this experiment, we select three domain-specific datasets from open-source CQA platforms: AskUbuntu, ServerFault, and Unix, all of which originate from a low-resource domain with minimal properly annotated data available on these topics. These datasets, considered from June 2023, includes questions (title and body), a list of answers, an accepted answer flag, tags for the questions, and the posting dates and times for both questions and answers. For each question, the accepted answer serves as the ground truth. Due to the limited resources of the datasets and the high expenses associated with human involvement, we opt not to use human annotations. Further we apply several filtering procedures, such as duplicate question removal, non-specific answer removal, and length constraints (token limit in LLM) resulting in our dataset. Adopting the temporal splitting approach inspired by Hazra et al. (2021, 2023), we consider training set from 2019-2020 and test dataset from 2021-2023. Due to resource limitations, we randomly sample test dataset, the details of which are provided in Table 1. From the training set of each dataset, we construct an instruction-tuning dataset by pairing questions and answers in the format ‘$[INST]Question[\textbackslash INST]Answer:actual\_answer$’. We prepare this instruction data for each dataset.
@@ 5 Experimental setup
Pre-LLM era baselines: We compare our approach with SOTA answer generation/summarization works such as AnswerBot Xu et al. (2017); Cai et al. (2019), GenQA Hsu et al. (2021) and TechSumBot Yang et al. (2023a) (see Appendix A.2). Due to unavailability of the codebase and unclear implementation details, we could not compare this paper Deng et al. (2019) with our method. Zeroshot LLMs: In this setting, we use various competitive LLMs to generate the answer of the given question. The LLMs are of different parameter sizes (7B to 40B). Such a choice enables us to understand how well models with diverse parameter sizes perform in zero shot setting. [w/o INST] TextGen: In this setup, we use a vector database (chromaDB [^9] https://docs.trychroma.com/getting-started and FAISS [^10] https://python.langchain.com/docs/integrations/vectorstores/faiss) containing all the training set questions. We compute contextual similarity between query $q$ and all the questions in database. We rank the questions in database based on the cosine similarity scores (higher scores get top ranks) and retrieve top $k$ questions. Further we use the top $k$ questions and their actual answers as few shot examples to the pretrained LLM for generating the answer. [w/o INST] TextContextGen: In this setup, we retrieve the top $k$ questions and their answers using the same method as  [w/o INST] TextGen. Subsequently, we use context enhancement component of our approach to enhance the context. Further we provide the enhanced context and the query as input to the pretrained LLM and obtain the generated answer. [w/o INST] GraphGen: In this setup, we use the Retriever module of our algorithm to retrieve the top $k$ questions. We use $k$ questions and their answers as few shot examples to the pretrained LLM for generating the answer. [w/o INST] GraphContextGen: We follow the retrieval step from  [w/o INST] GraphGen. Further we use our ContextEnhancer module to enhance the context. FineTuned Gen Zero-Shot: In this setting, we use instruction fine tuned LLM in zero shot settings. Here, we pass the questions as input and the fine tuned LLM generates the answer. TextGen: This setup is same as [w/o INST] TextGen. However, we use our instruction fine tuned LLM for generation. GraphGen: This setup is same as [w/o INST] GraphGen. However, we use our instruction fine tuned LLM for generation.
@@ 6 Results
Zero-shot answer generation by different LLMs: For AskUbuntu, Llama2 (13B) achieves the highest BERTScore (0.810), ROUGE 1 (0.227), and ROUGE L (0.211). For Unix, Llama2 (13B) leads in BERTScore (0.806), while Vicuna (13B) tops ROUGE 1 (0.181), and Llama2 (13B) tops ROUGE L (0.175). For ServerFault, Llama2 (13B) dominates BERTScore (0.819), Vicuna (13B) leads in ROUGE 1 (0.185), and Llama2 (13B) in ROUGE L (0.176). Performances are not always proportional to model size, as Llama2 (13B) often outperforms larger models like Gpt-neox (20B) and Falcon (40B). Models of similar sizes also display varied performances, indicating the importance of architecture and training methods.
Main results: Table 3 compares baseline results with our proposed method. Pre-LLM era: AnswerBot shows competitive BERTScore (0.803, 0.791, 0.802) for AskUbuntu, Unix, and ServerFault, respectively, while GenQA underperforms on Unix (BERTScore 0.55). AnswerBot achieves the highest FactSumm score across all platforms. LLM era: Llama2 (13B) is the reference model for generating answers. Our model, GraphContextGen, outperforms all baselines in BERTScore and FactSumm for AskUbuntu and Unix, producing factually more correct answers. For ROUGE 1 and ROUGE L, GraphContextGen shows competitive performance in Unix and ServerFault. Models from the LLM era generally outperform pre-LLM models in these metrics.
Grounding of the generated answers: We use UniversalNER Zhou et al. (2023) to identify entities in the ground truth and generated answers. The Jaccard similarity between entity sets for our model is 0.85, 0.75, and 0.79 for AskUbuntu, Unix, and ServerFault, respectively (Figure 2(A)). The overlap in the number of triplets is shown in Figure 2(B), indicating our model’s answers are rich in entities and relationships present in the ground truth.
@@
6.1 Ablation study
  • Based on embedding algorithms: The performance differences between sentence BERT Reimers and Gurevych (2019) and BGE embeddings Xiao et al. (2023) within the TextGen model are evident. The BGE variant demonstrates a slightly improved performance over sentence BERT.
  • Based on one-shot: When focusing on one-shot methodologies, the GraphGen model achieves marginally better scores compared to the TextGen model. This suggests that the incorporation of graph structures potentially aids in better knowledge retrieval
Method AskUbuntu Unix ServerFault
Based on embedding algorithms
[w/o INST] TextGen (Sentence BERT)
0.808 0.783 0.803
[w/o INST] TextGen (BGE)
0.812 0.809 0.810
Based on one-shot
[w/o INST] TextGen
0.809 0.807 0.807
[w/o INST] GraphGen
0.818 0.809 0.812
Based on only knowledge graph context
[w/o INST] OnlyContextGen[FT]
0.819 0.807 0.819
[w/o INST] OnlyContextGen[FG]
0.827 0.820 0.818
7 Human evaluation
In the results of the automatic evaluation, it is essential to remember that low values of BERTScore or FactSumm might also correspond to lower-quality ground truth answers posted by humans. Thus, there is a possibility that the model-generated answers are superior in quality compared to the ground truth answers. Such points can be verified only by human judgment experiments presented in Figure 3 and the Appendix A.6.
8 Error Analysis
This section presents a systematic error analysis highlighting the error types and corresponding examples (see Table 7).
9 Complexity analysis
[…]
10 Conclusion
11 Limitation
Using knowledge graphs with LLMs to generate technical answers encounters several challenges. The effectiveness of this approach partially depends on the accuracy and completeness of the knowledge graphs, as any gaps or inaccuracies can mislead the LLM’s responses. LLMs tailored for summarization may sacrifice detail for conciseness, potentially overlooking critical nuances of technical topics. Challenges in handling ambiguous contexts and the potential for biases introduced during knowledge graph construction and LLM training further complicate accurate answer generation. Moreover, the computational resources required for processing large-scale knowledge graphs pose scalability issues. Ensuring that knowledge graphs remain up-to-date and that LLMs can adapt to new information without frequent retraining are ongoing concerns. Finally, while LLMs are versatile, they may not achieve the level of specificity and accuracy provided by systems specialized in particular non niche domains.
12 Ethical consideration
The information in our dataset is free from harmful or offensive materials. We take serious measures to anonymize and handle any personal or sensitive data with the highest level of confidentiality. Protection of participants’ privacy is our priority and we consistently ensure to acquire their informed consent when collecting, annotating, and analyzing data. We provide equal incentives to all annotators for their efforts toward the annotation task.
Appendix A Appendix
A.1 Instruction tune hyperparameters
Hyperparameter Value
Learning Rate 2e-4
Batch Size 4
Gradient Accumulation Step 1
Number of Epochs 10
Weight Decay 0.001
Optimizer paged adamw 32bits
LR scheduler cosine
Warmup ratio 0.03
Max grad norm 0.3
bf16 True
LoRA r, alpha, dropout 64, 16, 0.1
bf16 True
Quantization 4bit
PEFT Techniques LoRA
Trainer SFTT
Question Related Question from PPR Related Question from simple similarity
Experiencing Intermittent Network Failures
on Ubuntu Server After Recent Update.
Desc: After a recent update on my Ubuntu
20.04 server, I’m experiencing intermittent
network failures. The server loses connectivity
randomly, and I’ve been unable to diagnose the
issue. Here’s the output of ifconfig and
dmesg | grep eth0 after the failure occurs…
How can I rollback a recent Ubuntu update to
troubleshoot network connectivity issues?
Desc: Following recent network issues on
my Ubuntu server, I suspect a recent update
might be the cause. I need to rollback this
update to confirm. What is the safest way to
revert the last system update? Is there a way
to identify which packages were updated
and selectively rollback, or do I need to restore
from a backup?
What are some common network troubleshooting
tools in Ubuntu for diagnosing connectivity problems?
Desc: In dealing with intermittent network failures on
my Ubuntu server, I’m looking for effective tools or
commands to diagnose the issue. What are the best tools
available in Ubuntu for network troubleshooting, especially
for monitoring and logging network activity over time to
catch these intermittent failures?
Script for Automating Log File Rotation
and Compression Not Working as Expected
on Linux.
Desc: I’m attempting to create a bash script
to automate log file rotation and compression
in a Linux environment. The script is supposed
to find all log files under /var/log , compress them,
and then move them to /archive/logs.However,
it’s not working as expected, and some log files
are being missed…
How can I set up a cron job to run this script
daily at midnight?
Desc: I have a script for log file rotation and
compression, but I’m not sure how to set it up
as a cron job to run automatically. What is the
correct way to schedule this script in cron to
run daily at midnight? Are there any specific
considerations for running such scripts as cron jobs?
What are the best practices for managing log files in
a Unix environment?
Desc: As I work on automating log file rotation and
compression, I want to ensure I’m following best
practices. What are the recommended strategies
for log file management in a Unix environment?
This includes considerations for log rotation frequency,
compression, archiving, and ensuring log integrity and security?
Table 6: Retrieved questions from PPR (graph structure based) and simple similarity for certain questions.